An employee is selected from a staff of 10 to supervise a certain project by selecting a tag at random from a box containing 10 tags numbered from 1 to 10. Find the formula for the probability distribution of X representing the number on the tag that is drawn.

Find the mean and variance of the random variable X


1
Expert's answer
2021-03-24T14:09:34-0400

An employee is selected from a staff of 10 to supervise a certain project by selecting a tag at random from a box containing 10 tags numbered from 1 to 10. Find the formula for the probability distribution of X representing the number on the tag that is drawn.

Find the mean and variance of the random variable X


Let random variable X represent the number of the drawn tag. Since the simple space consists of 10 equally likely events

"\Omega=\{1, 2, 3, 4 ,5 ,6 ,7 ,8 ,9 ,10\}"

thus the formula for probability distribution

"f(x)=\begin{cases} \frac{1}{10} \ for \ x=1, 2, ...,10\\ 0 \ elsewhere \end{cases}"


The mean:

"\mu=\sum x_i P(x_i)=0.1\cdot(1+2+3+4+5+6+7+8+9+10)=5.5"


Variance:

"var(x)=\sum(x_i-\mu)^2P(x_i)=0.1\cdot((1-5.5)^2+(2-5.5)^2+(3-5.5)^2+(4-5.5)^2+(5-5.5)^2+(6-5.5)^2+(7-5.5)^2+(8-5.5)^2+(9-5.5)^2+(10-5.5)^2)=0.1\cdot(20.25+12.25+6.25+2.25+0.25+0.25+2.25+6.25+12.25+20.25)=8.25"